Back to Main Archeops Page       Back to Main Journal Page      Back to KirunaS2.html

Daily reports of December 2001/January 2002

Local contact:      ESRANGE, Po Box 802, 3981-28 Kiruna, Suède   Tel  46 980 72000,  Fax 46 980 72 254
Archeops balloon Hall telephone number: +46 980 72 267

Latest news from the Archeops team

Tuesday 20th, November 2001
Shipping by lorry for Kiruna. Expected arrival on Monday 26th.

Monday 26th, November 2001
Alain, Karine, Xavier, Jean-Charles, Jean-Christophe and Nicolas arrived yesterday evening. We unloaded the truck this morning and set everything up imediately. Everything is ok, nothing is broken on the cryostat nor on the polarizers matrix. Alain and Karine are closing the cryostat tonight ! Armando and Federico are arriving this evening. The meteo man is arriving tonight as well and we'll therefore have our first meteo meeting on wednesday.

Tuesday 27th, November 2001
We started to cool down the cryostat and everything is fine. The meteo chief is now at the base and will give us the first news tomorrow. Armando and Ferico set up the stellar sensor. We all spent some time on the software, but Alain also pointed out a new list of things to do on the hardware, like setting the crash pads, do some baffling... !

Wednesday 28th, November 2001
 Yesterday evening we picked up Michel Piat, Cyrille Rosset and Bertrand Gautier at the airport who brought with him the last pieces of electronics hardware. The cryostat is behaving very well but plug in the transfert tube. We'll warm up and restart tomorrow, taking advantage of this delay to finalize the flight configuration in the flight EPROM. The CNES opening flight of the campaign is scheduled for next monday if the weather gets better. For hist first day in winter in Kiruna, Cyrille experienced the installation of the thermal source with Jean-Christophe and Nicolas on Radar Hill, and seemed to enjoy it. Armando and Federico mounted the stellar sensor on the gondola, then wrapped it with its shields.

Thursday 29th, November 2001
The mirrors were cleaned and some more baffling was done. The last touch on the acquisition software was implemented to enable the reading of gyroscopes data in Archediab. The cryostat is being cooled down. The inmarsat antenna is being mounted.

Friday 30th, November 2001
Jean-Charles and Bertrand finished the flight batteries. Armando and Federico did great job on the cryostat baffling. Meanwhile, all the others were swithing from hard- to software. The cryostat is cooling down and is in good health !

Saturday 1st, December 2001
We had this morning the first meeting with the CNES to open the balloon campaign ! The flight chain and telemetry were defined. The wind conditions in the stratosphere are bad at the moment but could get better around the end of next week, so we are still actively preparing everything. As we had to buy some food, we took the opportunity to take some pictures in Kiruna. The cryostat is now at 133mk and will be ready for photometry measurements tomorrow morning.

Sunday 2nd, December 2001
One more member of the team arrived today ! Xavier picked up Christophe Magneville at 7pm. The cryostat had cooled down to 85mK with a bolometer plate at 95mK this morning and we performed the photometry calibration. It appears that the lenses we added at the last moment damp the optical efficiency. We therefore decided to warm up the cryostat, open it up and turn back to the previous Grenoble configuration.

Monday 3rd, December 2001
This morning we opened up the cryostat and removed the filters and lenses. The cryostat was then closed back in the mid afternoon and is now cooling down perfectly. The inmarsat antenna works fine too. Meanwhile we saw two ballon launch : the first one of Ozone Sampler a German experiment, and the second one of CNES which was testing a new prototype of Presurized Balloons. The weather forecast are confirmed : the campaign opening flight is scheduled tomorrow, but the vortex is not in place yet. The possibility of a flight for Archeops by the end of next week remains... and if you need some distraction after this tiring sunday ...

Tuesday 4th, December 2001
We mounted the cryostat on the gondola. It is cooling down perfectly and is 1.7K thanks to Karine, Alain and their night shift. We started to calibrate the gyroscopes and to equilibrate the weight distribution of the gondola. Beside this, the opening flight of CNES this morning failed at take off due to a leak in the balloon. They will try tomorrow again. Our friends of MIPAS arrived last night are now fully set up right next to us ! (+ see 2001/12/01 for 2 new nice pictures)

Wednesday 5th, December 2001 We tested the CNES telemetry this morning and it worked perfectly. The calibration of gyroscopes is on its way. We also performed some pointing tests : the stellar sensor is focused, and the maps on the thermal source will soon be reduced. Jean arrived this evening at 23.30. The CNES succesfully launched the opening campaign balloon at 17.30 and will launch a second one of 25000 cubic meters tomorrow morning.

Thursday 6th, December 2001 (Sunrise etc) (what would have been the trajectory)

- Tests of gyros and magnetometer by rotating the gondola inside the hall. Magnetometer and gyros work OK and are calibrated.
- Calibration cryostat mounted and calibration performed. Analysed results tomorrow.
- Anniversaire de Karine (26 ans !!!) - with many very very very nice pictures.

Friday 7th, December 2001

- Frederico, Armando and Nicolas departure at 4.30 am.
- Baffling (re-)mounted.
- New calibration performed with calibration cryostat. Today and yesterday calibrations are consistent. Compare with previous flight the 143GHz are much more efficient and the 353 Ghz are identical. The 217 Ghz are working well and we will get 7 bolometers for that flight. Unfortunatly, we will have only one bolometer at 545 Ghz. Optical calibration waiting next pointing to the thermal source.
- Test of the synchronous source: laser and thermal source synchros are working. There is a problem of bad connection inside the antenna to be solved.
- a marvelous Northern Light to close this day.

Saturday 8th, December 2001

- Cryostat (re-)mounted on the gondola.
- Pointing to Gunn diode and thermal source with synchro. 143 Ghz bolo have about 8 arcmin fwhm. 217 Ghz are all responding. Tests 353 Ghz with polarised thermal source.
- Juan has arrived, Jean-Christophe and Cyrille have left this morning.

Sunday 9th, December 2001

We are on stand-by till thursday for more favorable wind conditions. Today we looked at yesterday data and the results make sens. Alain contined improving the quick look analysis. We had our first day without clouds.

Monday 10th, December 2001

- PSC experiment succesfully launched yesterday night around 9 pm.
- The vortex position is still not good, even worse. At present the trajectory goes to New Zemble island ! Everyone is waiting for the same thrilling experiment...
- Jean-Charles, Michel and Bertand have left this morning.
- Sofware developement for raw map making during the flight under improvement.
- Baffling is now finished and ready for flight.
- Map with gaussian fits for main lobes, following the calibration made last friday.

Tuesday 11th, December 2001

- Vortex positon and wind direction are not good, even worst than yesterday. Next possible launch window around Monday. Still waiting!
- Gondola equilibrium performed.

Wednesday 12th, December 2001

- Still waiting for good stratospheric winds. (a patience game to forget the vortex ...)
- Alexandre and Karine gave a talk on Archeops in Kiruna University.
- Quick look analysis programs for ultra compressed data of all bolometers in progress.
- Jacques just arrived in Moscow.
- First time that Archeops see a distant (1.e-13 light year = 1km!!!) polarised signal. Verification of the polarisers angle.

Thursday 13th, December 2001

- Testing Inmarsat link with gondola outside.
- A lot of wind on the launch area today.
- Still bad meteo for stratospheric winds.

Friday 14th, December 2001

Still waiting !

Saturday 15th, December 2001 (12-16)

- Kiruna sight seeing for most of us. On return came the bad news:
Due to bad meteo conditions for stratospheric wind, it will be impossible to launch before December the 22th.

- Alain, Xavier, Juan, Karine, Jean and Christophe will leave tomorrow on Sunday December the 16th.
- Alexandre and Jean-philippe will leave on Monday December the 17th.

Predicted trajectories (Dec 2001)    Celestial infomation (Dec 2001)

Kiruna 2001/2002 Home Page